OpenStudy (yinshy):

Use an addition property to solve for c. 14 + (–12) = –12 + c A. –26 B. –14 C. 14 D. 26

OpenStudy (yinshy):

@Awolflover1

OpenStudy (yinshy):

@shaima

Awolflover1 (awolflover1):

\[14−12=−12+c\]

Awolflover1 (awolflover1):

\[2=c−12\]

Awolflover1 (awolflover1):

now flip it \[c−12=2\]

OpenStudy (yinshy):

is it C???

Awolflover1 (awolflover1):

yes

Awolflover1 (awolflover1):

\[c−12+12=2+12\]

OpenStudy (yinshy):

derp

OpenStudy (yinshy):

to EZ

OpenStudy (yinshy):

it gives you the Awnser

OpenStudy (yinshy):

xD